# Probing Reheating Phase via Non-Helical Magnetogenesis and Secondary Gravitational Waves

### v1

@@ Abstract
In the past two decades, significant advancements have been made in observational techniques to enhance our understanding of the universe and its evolutionary processes. However, our knowledge of the post-inflation reheating phase remains limited due to its small-scale dynamics. Traditional observations, such as those of the Cosmic Microwave Background (CMB), primarily provide insights into large-scale dynamics, making it challenging to glean information about the reheating era. In this paper, our primary aim is to explore how the generation of Gravitational Waves (GWs) spectra, resulting from electromagnetic fields in the early universe, can offer valuable insights into the Reheating dynamics. We investigate how the spectral shape of GWs varies across different frequency ranges, depending on the initial magnetic profile and reheating dynamics. For this, we consider a well-known non-helical magnetogenesis model, where the usual electromagnetic kinetic term is coupled with a background scalar. Notably, for such a scenario, we observe distinct spectral shapes with sufficiently high amplitudes for different reheating histories with the equation of state parametrized by ($w_{\rm re}$). We identify spectral breaks in the GW spectra for both $w_{\rm re}<1/3$ and $w_{\rm re}>1/3$ scenarios. We find that future GW experiments such as BBO, LISA, SKA, and DECIGO are well within the reach of observing those distinct spectral shapes and can potentially shed light on the underlying mechanism of the reheating phase. Finally, we attempt to constrain the non-helical models under consideration utilizing the latest NANOGrav 15-year observation campaign. We find that in order to explain the observed GW spectrum in the nano-Hz range, the required magnetic field spectrum needs to be very stiff. However, such a stiff magnetic field spectrum in turn yields a strongly blue-tilted GW spectrum, which is in conflict with other cosmological observations.
@@ I Introduction:
Studies on inflationary magnetogenesis Kandus et al. (2011); Turner and Widrow (1988); Durrer and Neronov (2013); Ferreira et al. (2013); Subramanian (2016); Kobayashi (2014); Haque et al. (2021b); Tripathy et al. (2022); Li and Zhang (2022); Adshead et al. (2016); Campanelli (2009); Jain et al. (2014); Caprini and Sorbo (2014); Bamba et al. (2022); Maiti et al. (2025) has mostly focused on producing large-scale magnetic field (LSMF) in the intergalactic medium in voids ($\sim 10^{-16}$ Gauss of coherence length as large as Mpc scales) as hinted by different observations Neronov and Vovk (2010); Essey et al. (2011); Saga et al. (2018). Furthermore, such large-scale magnetic fields are also argued to act as a tiny seed initial magnetic field which may be amplified by the well-known galactic dynamo mechanism Fletcher (2011); Beck (2016); Haverkorn et al. (2008); Kronberg et al. (2001) to produce magnetic fields at various astrophysical and galactic scales of order a few micro Gauss Grasso and Rubinstein (2001); Kronberg et al. (2001); Turner and Widrow (1988). Non-linear magneto-hydrodynamic effects at astrophysical and galactic scales generically erase the detailed properties of the initial seed field, which makes it difficult to verify such inflationary mechanisms of magnetogenesis models. Moreover, magnetic fields in the voids of the Mpc scale are extremely challenging to detect. Apart from the indirect $\gamma$-ray observational lower limit of LSMF setting some constraints Acciari et al. (2023); Takahashi et al. (2011); Arlen et al. (2014), all the magnetogenesis models are largely unconstrained and poorly understood from the observational point of view. Therefore, it is essential to look for some indirect observables that can help understand such mechanisms better. In this realm, studies have been performed on the physical effect of LSMF on various important cosmological observables such as inflationary power spectrum Hortúa and Castañeda (2014), CMB Paoletti et al. (2022a); Zucca et al. (2017), gravitational waves Sorbo (2011); Caprini and Durrer (2001); Ito and Soda (2017); Sharma et al. (2020); Okano and Fujita (2021). In this paper, we shall study in detail the production of SGW due to the electromagnetic field produced during inflation. As we shall observe, the SGW dynamics and its spectrum crucially depend on the magnetogenesis model and the background evolution. Therefore, SGW naturally encodes the physics of inflation and the magnetogenesis models together.
@@
After the end of inflation Guth (1981); Linde (1982); Albrecht and Steinhardt (1982); Starobinsky (1980); Liddle et al. (1994), both GWs and the produced electromagnetic field will evolve through a reheating phase which depends on the very nature of the inflaton field near its minimum and how it is coupled to the standard model fields. In the simplest perturbative framework, this phase is parametrized by the reheating temperature $(T_{\rm re})$ which encodes inflaton decay width, $\Gamma$ into standard model fields through the relation $T_{\rm re}\propto\sqrt{\Gamma}$, and the inflation equation of state ($w_{re}$) which encodes the nature of the inflaton potential Kolb and Turner (1990); Kofman et al. (1997); Bassett et al. (2006); Allahverdi et al. (2010); Amin et al. (2014); Drewes et al. (2017); Haque and Maity (2023a). Even though the reheating temperature is broadly bounded within $10^{15}\mbox{GeV}>T_{\rm re}>T_{BBN}\sim 10~{}\mbox{MeV}$ Cook et al. (2015); Haque and Maity (2023b), the inflaton/reheating equation of state is mostly unconstrained (see recent works on this Haque and Maity (2023b)). A recent surge of activity on understanding this phase has been to realize the underlying connection among inflationary parameters such as scalar spectral index $(n_{s})$ and the reheating parameters $(T_{\rm re},w_{re})$Dai et al. (2014); Martin and Ringeval (2010); Maity and Saha (2018); Haque and Maity (2023b). Furthermore, the evolution of PGW has also been shown to encode valuable information when passing through this phase, see Haque et al. (2021a) and references therein. In this paper, we will analyze in detail the production of SGW, a specific class of non-helical magnetogenesis scenario accounting for the observable constraints on the present-day LSMF combined with the CMB anisotropies, BBN, and other sensitivity curves of the future observatories. Our SGW analysis is observed to put interesting constraints on the nature of inflation through its equation of state $w_{re}$, reheating temperature $T_{\rm re}$, and the inflationary magnetogenesis model itself. To generate the LSMF, we consider a simple model of primordial magnetogenesis with the coupling $f^{2}(\phi)FF$ Subramanian (2016); Kobayashi (2014); Haque et al. (2021b); Tripathy et al. (2022).
@@ II Magnetogenesis: quantizing the gauge field and evolution of its power spectrum
To make our paper self-contained, we briefly describe the quantization of gauge fields and define the associated spectrum in the FLRW background. Maxwell’s theory in four dimensions is conformally invariant, which means that to generate magnetic fields in a spatially flat universe as discussed in Kobayashi (2014); Ferreira et al. (2013); Subramanian (2016); Haque et al. (2021b); Tripathy et al. (2022); Campanelli (2009); Jain et al. (2014); Caprini and Sorbo (2014); Bamba et al. (2022), one must break conformal invariance. To achieve this, a time-dependent effective gauge coupling $f(\eta)$ is typically introduced Kobayashi (2014); Ferreira et al. (2013); Subramanian (2016); Haque et al. (2021b); Tripathy et al. (2022).
@@
where ‘$\eta$’ is the conformal time. Spatial flatness helps us to denote the vector potential in terms of irreducible components as $A_{\mu}=\left(A_{0},\partial_{i}S+A_{i}\right)$ with the traceless condition $\delta^{ij}\partial_{i}A_{j}=0.$ In terms of these components the action Eq.(1) becomes,
[…]
@@ II.1 Defining Reheating Temperature and Inflationary Parameters
where $\alpha={1}/{3(1+w_{re})}$ and $T_{0}=2.725\,\text{K}$ is the present-day CMB temperature. Here, we consider $a_{0}=1$ as the present-day value of the scale factor.
@@ II.2.1 Evolution During inflation:
$f(\eta)=\left\{\begin{matrix}\left(\frac{a}{a_{end}}\right)^{n}&\mbox{ for $a\leq a_{end}$}\\ 1&\mbox{ for $a>a_{end}$}\end{matrix}\right..$ (11)
Note that after the end of inflation set at $a_{end}$, the coupling function becomes unity, which restores the usual conformal symmetry of the electromagnetic field.
@@
$c_{1}=c_{3}=0~{};~{}c_{2}=\frac{k^{4}}{8\pi}\left(\frac{1}{-k\eta_{end}}\right)^{2n}.$ (13)
@@
After obtaining the power spectrum during inflation, we can use it as a boundary condition to solve for the power spectrum during reheating. It is important to note that for $n>1/2$, the ration of two strength turned out to be $\mathcal{P}_{B}^{\mathrm{inf}}(k)/\mathcal{P}_{E}^{\mathrm{inf}}(k)\propto(k\eta)^{-2}>>0$. This immediately suggests that at the end of inflation, the magnetic (electric) field strength dominates over the electric (magnetic) field strength for $n>1/2(<-1/2)$. To estimate the magnetic field strength at the end of inflation for a specific set of parameters, we consider an inflationary energy scale of $H_{{}_{\mathrm{I}}}\simeq 10^{-5}M_{\rm P}$ and an inflationary duration corresponding to $N_{{}_{\mathrm{I}}}=55$ e-folds. For $n=2$, which corresponds to a scale-invariant magnetic spectrum, the field strength at a comoving scale of $k=1~{}\mathrm{Mpc}^{-1}$ is found to be $B_{n=2}(1~{}\text{Mpc}^{-1})\simeq 10^{46}$ G. In contrast, for $n=-2$, where the magnetic spectrum exhibits a scale-dependent behavior with $B_{n=-2}\propto k$, the field strength at the same scale is estimated as $B_{n=-2}(1~{}\mathrm{Mpc}^{-1})\simeq 10^{23}$ G. Notably, for scale-dependent spectra, the field strength is influenced by the duration of inflation due to the factor $(k/k_{\rm e})^{n_{B}}$, whereas for scale-invariant spectra, it remains independent of the inflationary duration.
@@
$B_{0}(k)\simeq 1.84\times 10^{-12}\mbox{ G}\left(\frac{k}{k_{\rm e}}\right)^{2-|n|}\times\left\{\begin{matrix}2^{n+\frac{1}{2}}\Gamma(n+\frac{1}{2})&\mbox{ for $n>1/2$}\\ 2^{n-\frac{1}{2}}\Gamma(n-\frac{1}{2})\frac{k}{k_{\rm e}}&\mbox{ for $n<-1/2$}\end{matrix}\right..$ (18)
@@
Notably, the case of $n=2$ yields a scale-invariant magnetic field. Despite the scale-invariance, it faces the challenge of the strong coupling problem due to the large effective electromagnetic gauge coupling $\propto 1/f^{2}$, at the onset of inflation. However, it can generate a magnetic field strength of around $\sim 10^{-11}$ G. Note that such a strong magnetic field can have a sizable impact on the tensor-to-scalar ratio, which we will discuss in the GW section. On the other hand, the $n=-2$ coupling leads to a scale-dependent magnetic field strength is $B_{0}(1\,\mathrm{Mpc}^{-1})\simeq 1.45\times 10^{-36}$ G, which is very small compared with the current observational bound on the 1 Mpc scale Neronov and Vovk (2010); Essey et al. (2011); Saga et al. (2018).
Backreaction of Electromagnetic Fields:
Including the coupling function $f(\eta)$ in Eq. (1) facilitates electromagnetic (EM) field generation during inflation, resulting in a total post-inflation energy density $\rho_{\mathrm{total}}=\rho_{\mathrm{c}}^{\mathrm{I}}+\rho_{\mathrm{em}}^{\mathrm{I}}$, where $\rho_{\mathrm{c}}^{\mathrm{I}}=3H_{{}_{\mathrm{I}}}^{2}M_{\rm P}^{2}$ is the inflaton energy density and electromagnetic energy density

### v2

@@ Abstract
In the past two decades, significant advancements have been made in observational techniques to enhance our understanding of the universe and its evolutionary processes. However, our knowledge of the post-inflation reheating phase remains limited due to its small-scale dynamics. Traditional observations, such as those of the Cosmic Microwave Background (CMB), primarily provide insights into large-scale dynamics, making it challenging to glean information about the reheating era. In this paper, our primary aim is to explore how the generation of Gravitational Waves (GWs) spectra, resulting from electromagnetic fields generated during inflation, can offer valuable insights both into the magnetogenesis model and Reheating dynamics. We investigate how the spectral shape of GWs varies across different frequency ranges, depending on the model parameters, the magnetogenesis parameters, and reheating dynamics. For this, we consider a well-known non-helical magnetogenesis model, where the usual electromagnetic kinetic term is coupled with a background scalar. Notably, for such a scenario, we observe distinct spectral shapes with sufficiently high amplitudes for different reheating histories with the equation of state parametrized by ($w_{\rm re}$). We identify spectral breaks in the GW spectra for both $w_{\rm re}<1/3$ and $w_{\rm re}>1/3$ scenarios. We find that future GW experiments such as BBO, LISA, SKA, and DECIGO are well within the reach of observing those distinct spectral shapes and can potentially shed light on the underlying mechanism of the reheating phase. Finally, we attempt to constrain the non-helical models under consideration utilizing the latest NANOGrav 15-year observation campaign. We find that in order to explain the observed GW spectrum in the nano-Hz range, the required magnetic field spectrum needs to be very stiff. However, such a stiff magnetic field spectrum in turn yields a strongly blue-tilted GW spectrum, which conflicts with other cosmological observations.
@@ I Introduction:
Studies on inflationary magnetogenesis Ratra (1992); Demozzi et al. (2009); Kandus et al. (2011); Turner and Widrow (1988); Durrer and Neronov (2013); Ferreira et al. (2013); Subramanian (2016); Kobayashi (2014); Haque et al. (2021b); Tripathy et al. (2022); Li and Zhang (2022); Adshead et al. (2016); Campanelli (2009); Jain et al. (2014); Caprini and Sorbo (2014); Bamba et al. (2022); Maiti et al. (2025); Papanikolaou et al. (2025) has mostly focused on producing large-scale magnetic field (LSMF) in the intergalactic medium in voids ($\sim 10^{-16}$ Gauss of coherence length as large as Mpc scales) as hinted by different observations Neronov and Vovk (2010); Essey et al. (2011); Saga et al. (2018). Furthermore, such large-scale magnetic fields are also argued to act as a tiny seed initial magnetic field which may be amplified by the well-known galactic dynamo mechanism Fletcher (2010); Beck (2016); Haverkorn et al. (2008); Kronberg et al. (2001) to produce magnetic fields at various astrophysical and galactic scales of order a few micro Gauss Grasso and Rubinstein (2001); Kronberg et al. (2001); Turner and Widrow (1988). Non-linear magneto-hydrodynamic effects at astrophysical and galactic scales generically erase the detailed properties of the initial seed field, which makes it difficult to verify such inflationary mechanisms of magnetogenesis models. Moreover, magnetic fields in the voids of the Mpc scale are extremely challenging to detect. Apart from the indirect $\gamma$-ray observational lower limit of LSMF setting some constraints Acciari et al. (2023); Takahashi et al. (2011); Arlen et al. (2014), all the magnetogenesis models are largely unconstrained and poorly understood from the observational point of view. Therefore, it is essential to look for some indirect observables that can help understand such mechanisms better. In this realm, studies have been performed on the physical effect of LSMF on various important cosmological observables such as inflationary power spectrum Hortúa and Castañeda (2014), CMB Paoletti et al. (2022); Zucca et al. (2017), gravitational waves Sorbo (2011); Caprini and Durrer (2001); Ito and Soda (2017); Sharma et al. (2020); Okano and Fujita (2021). In this paper, we shall study in detail the production of SGW due to the electromagnetic field produced during inflation. As we shall observe, the SGW dynamics and its spectrum crucially depend on the magnetogenesis model and the background evolution. Therefore, SGW naturally encodes the physics of inflation and the magnetogenesis models together.
@@
After the end of inflation Guth (1981); Linde (1982); Albrecht and Steinhardt (1982); Starobinsky (1980); Liddle et al. (1994), both GWs and the produced electromagnetic field will evolve through a reheating phase which depends on the very nature of the inflaton field near its minimum and how it is coupled to the standard model fields. In the simplest perturbative framework, this phase is parametrized by the reheating temperature $(T_{\rm re})$ which encodes inflaton decay width, $\Gamma$ into standard model fields through the relation $T_{\rm re}\propto\sqrt{\Gamma}$, and the inflation equation of state ($w_{re}$) which encodes the nature of the inflaton potential Kolb and Turner (1990); Kofman et al. (1997); Bassett et al. (2006); Allahverdi et al. (2010); Amin et al. (2014); Drewes et al. (2017); Haque and Maity (2023a); Adhikari et al. (2022); Yogesh et al. (2024); Ghosh and Ghosh (2025); Gangopadhyay et al. (2023); Khan and Yogesh (2022). Even though the reheating temperature is broadly bounded within $10^{15}\mbox{GeV}>T_{\rm re}>T_{BBN}\sim 10~\mbox{MeV}$ Kawasaki et al. (1999); Cook et al. (2015); Haque and Maity (2023b); Hasegawa et al. (2019), the inflaton/reheating equation of state is mostly unconstrained (see recent works on this Haque and Maity (2023b)). A recent surge of activity on understanding this phase has been to realize the underlying connection among inflationary parameters such as scalar spectral index $(n_{s})$ and the reheating parameters $(T_{\rm re},w_{re})$Dai et al. (2014); Martin and Ringeval (2010); Maity and Saha (2018); Haque and Maity (2023b); Adhikari et al. (2022); Yogesh et al. (2024, 2025). Furthermore, the evolution of PGW has also been shown to encode valuable information when passing through this phase, see Haque et al. (2021a) and references therein. In this paper, we will analyze in detail the production of SGW, a specific class of non-helical magnetogenesis scenario accounting for the observable constraints on the present-day LSMF combined with the CMB anisotropies, BBN, and other sensitivity curves of the future observatories. Our SGW analysis is observed to put interesting constraints on the nature of inflation through its equation of state $w_{re}$, reheating temperature $T_{\rm re}$, and the inflationary magnetogenesis model itself. To generate the LSMF, we consider a simple model of primordial magnetogenesis with the coupling $f^{2}(\phi)FF$ Subramanian (2016); Kobayashi (2014); Haque et al. (2021b); Tripathy et al. (2022); Dimopoulos et al. (2025).
@@ II Magnetogenesis: quantizing the gauge field and evolution of its power spectrum
To make our paper self-contained, we briefly describe the quantization of gauge fields and define the associated spectrum in the Friedmann–Lemaître–Robertson–Walker (FLRW) background. Maxwell’s theory in four dimensions is conformally invariant, which means that to generate magnetic fields in a spatially flat universe as discussed in Kobayashi (2014); Ferreira et al. (2013); Subramanian (2016); Haque et al. (2021b); Tripathy et al. (2022); Campanelli (2009); Jain et al. (2014); Caprini and Sorbo (2014); Bamba et al. (2022), one must break conformal invariance. To achieve this, a time-dependent effective gauge coupling $f(\eta)$ is typically introduced Kobayashi (2014); Ferreira et al. (2013); Subramanian (2016); Haque et al. (2021b); Tripathy et al. (2022).
@@
where $\eta$’ is the conformal time. Spatial flatness helps us to denote the vector potential in terms of irreducible components as $A_{\mu}=\left(A_{0},\partial_{i}S+A_{i}\right)$ with the traceless condition $\delta^{ij}\partial_{i}A_{j}=0.$ In terms of these components the action Eq.(1) becomes,
$\mathcal{S}_{\rm EM}=\frac{1}{2}\int d\eta d^{3}xf(\eta)^{2}(A_{i}^{\prime}{A^{i}}^{\prime}-\partial_{i}A_{j}\partial^{i}A^{j}).$ (3)
@@ II.1 Defining Reheating Temperature and Inflationary Parameters
Here $k_{\rm e}$ is the highest mode that can leave the horizon at the end of inflation. In the above we have defined $\alpha={1}/{3(1+w_{re})}$ and $T_{0}=2.725\,\text{K}$ is the present-day CMB temperature. Here, we consider $a_{0}=1$ as the present-day value of the scale factor.
@@ II.2.1 Evolution During inflation:
$f(\eta)=\left\{\begin{matrix}\left(\frac{a}{a_{\rm end}}\right)^{n}&\mbox{ for $a\leq a_{\rm end}$}\\ 1&\mbox{ for $a>a_{\rm end}$}\end{matrix}\right.$. (11)
Note that after the end of inflation set at $a_{\rm end}$, the coupling function becomes unity, which restores the usual conformal symmetry of the electromagnetic field.
@@
$c_{1}=c_{3}=0~;~c_{2}=\frac{k^{4}}{8\pi}\left(\frac{1}{-k\eta_{\rm e}}\right)^{2n}.$ (13)
@@
After obtaining the power spectrum during inflation, we can use it as a boundary condition to solve for the power spectrum during reheating. It is important to note that for $n>1/2$, the ratio of the two field strengths turns out to be ${\mathcal{P}_{B}^{\mathrm{inf}}(k)}/{\mathcal{P}_{E}^{\mathrm{inf}}(k)}\propto(k\eta)^{-2}\gg 0$. This immediately suggests that, at the end of inflation, the magnetic field strength dominates over the electric field strength for $n>1/2$. However, it should be emphasized that the case $n>0$ suffers from a strong coupling problem. On the other hand, for $n<0$, the situation is reversed: the electric field strength dominates over the magnetic field strength, i.e. ${\mathcal{P}_{E}^{\mathrm{inf}}(k)}/{\mathcal{P}_{B}^{\mathrm{inf}}(k)}\propto(k\eta)^{-2}\gg 0$ and in this case there is no strong coupling issue. To estimate the magnetic field strength at the end of inflation for a specific set of parameters, we consider an inflationary energy scale of $H_{{}_{\mathrm{I}}}\simeq 10^{-5}M_{\rm P}$ and an inflationary duration corresponding to $N_{{}_{\mathrm{I}}}=55$ e-folds.Now we can define the magnetic field strength at the end of inflation as $B(k,\eta_{\rm e})=\sqrt{\mathcal{P}_{B}^{\mathrm{inf}}(k,\eta_{\rm e})}$. For $n=2$, which corresponds to a scale-invariant magnetic spectrum, the field strength at a comoving scale of $k=1~\mathrm{Mpc}^{-1}$ is found to be $B_{n=2}(1~\text{Mpc}^{-1})\simeq 10^{46}$ G. In contrast, for n=2n=-2, where the magnetic spectrum exhibits a scale-dependent behavior with $B_{n=-2}\propto k$, the field strength at the same scale is estimated as $B_{n=-2}(1~\mathrm{Mpc}^{-1})\simeq 10^{23}$ G. Notably, for scale-dependent spectra, the field strength is influenced by the duration of inflation due to the factor $(k/k_{\rm e})^{n_{B}}$, whereas for scale-invariant spectra, it remains independent of the inflationary duration.
@@
$B_{0}(k)\simeq 1.84\times 10^{-12}\mbox{ G}\left(\frac{k}{k_{\rm end}}\right)^{2-|n|}\times\left\{\begin{matrix}2^{n+\frac{1}{2}}\Gamma(n+\frac{1}{2})&\mbox{ for $n>1/2$}\\ 2^{n-\frac{1}{2}}\Gamma(n-\frac{1}{2})\frac{k}{k_{\rm end}}&\mbox{ for $n<-1/2$}\end{matrix}\right.$. (18)
@@
Notably, the case of $n=2$ yields a scale-invariant magnetic field. Despite the scale-invariance, it faces the challenge of the strong coupling problem due to the large effective electromagnetic gauge coupling $\propto 1/f^{2}$, at the onset of inflation. However, it can generate a magnetic field strength of around $\sim 10^{-11}$ G. Note that such a strong magnetic field can have a sizable impact on the tensor-to-scalar ratio, which we will discuss in the GW section. On the other hand, the n=2n=-2 coupling leads to a scale-dependent magnetic field strength is $B_{0}(1\,\mathrm{Mpc}^{-1})\simeq 1.45\times 10^{-36}$ G, which is very small compared with the current observational bound on the 1 Mpc scale Neronov and Vovk (2010); Essey et al. (2011); Saga et al. (2018).
II.2.2 Backreaction and Strong Coupling Problem
In models featuring a coupling of the form $f^{2}F_{\mu\nu}F^{\mu\nu}$, the effective electromagnetic gauge coupling is given by $\alpha_{\rm em}=e^{2}/f^{2}$. Consequently, if the coupling function satisfies $f\ll 1$ at the beginning of inflation, the effective gauge coupling becomes very large. This places the system in the strong coupling regime, rendering the perturbative treatment of the gauge field unreliable. Conversely, if the scenario is initialized with $f\gg 1$ and evolves toward $f=1$ by the end of inflation, thus restoring conformal symmetry, the effective gauge coupling remains small throughout, with $\alpha_{\rm em}\lesssim e^{2}$. In such cases, the model avoids the strong coupling problem.
For the specific choice of the coupling function discussed in Eq. (11), we find that for $n>0$, the initial value of the function is very small, $f(\eta_{i})\ll 1$, where $\eta_{i}$ denotes the beginning of the inflationary era. This implies that models with $n>0$ generally suffer from strong coupling issues at early times. In contrast, for $n<0$, the coupling function starts with a large value, avoiding the strong coupling problem altogether. As discussed in Eq. (17), for the regime n<1/2n<-1/2, the electric and magnetic power spectra exhibit the following behavior
$\mathcal{P}_{B}^{\mathrm{inf}}(k,\eta)=\frac{H_{{}_{\mathrm{I}}}^{4}}{8\pi}\frac{\Gamma^{2}\left(|n|-\frac{1}{2}\right)2^{2|n|-1}}{\pi^{2}}(-k\eta)^{n_{\mathrm{E}}+2},$ (19)
$\mathcal{P}_{E}^{\mathrm{inf}}(k,\eta)=\frac{H_{{}_{\mathrm{I}}}^{4}}{8\pi}\frac{2^{2|n|+1}\Gamma^{2}\left(|n|+\frac{1}{2}\right)}{\pi^{2}}(-k\eta)^{n_{\mathrm{E}}},$ (20)
where $n_{\mathrm{E}}=4-2|n|$ denotes the spectral index of the electric field. From the above discussion, it is evident that a scale-invariant magnetic power spectrum requires n=3n=-3. For this specific choice of the coupling parameter, the electric spectrum scales as $\mathcal{P}_{E}(k,\eta)\propto(-k\eta)^{-2}$, implying that in the limit $|-k\eta|\to 0$, the energy density of the electric field diverges. Consequently, the electric energy density may exceed the inflationary background energy density by the end of inflation. Therefore, the choice n=3n=-3 leads to a significant backreaction problem.
This consideration imposes both upper and lower bounds on the coupling parameter $n$. Beyond a certain threshold, the backreaction becomes negligible and the model remains consistent. To determine the lower bound on $n$, one must compute the total energy density associated with the electromagnetic field at the end of inflation, i.e., at conformal time $\eta=\eta_{\rm e}$. For $n<0$, the total energy density of the electromagnetic field is given by